### v2

# MR-STGN: Multi-Residual Spatio Temporal Graph Network using Attention Fusion for Patient Action Assessment

Abstract
Assessing patient actions is critical in healthcare because it aids in disease progression and treatment effectiveness. Traditional methods for assessing patient action, rely on manual observation and scoring, which can be subjective and time-consuming. In this paper, we propose a novel approach for patient action assessment using a Multi-Residual Spatio Temporal Graph Network (MR-STGN) that incorporates both angular and positional 3D skeletons. The MR-STGN is designed to capture the spatio-temporal dynamics of patient actions by integrating information from multiple residual layers, each of which extracts features at different levels of abstraction. To enhance the effectiveness of the MR-STGN, we incorporate an attention fusion mechanism that selectively combines the information from the angular and positional skeletons. This attention mechanism allows the network to adaptively weight the importance of different features, enabling it to focus on the most relevant aspects of the patient’s movements and give clear guidance on which body parts or movements to focus on. We evaluate our method in the UI-PRMD dataset, demonstrating that it outperforms state-of-the-art methods for automated patient action assessment. Our results show that the MR-STGN model can accurately predict patient actions score in real-time, with high precision and optimized time. We also conduct ablation studies to analyze the impact of different components of the proposed model.
Index Terms:
Automatic assessment, Rehabilitation, Spatio-Temporal, Graph convolution networks, Attention mechanism
I Introduction
Physical rehabilitation is a critical component of patient care, particularly for individuals with disabilities or injuries  [jang2019pulmonary]. However, ensuring patients are performing exercises correctly and adhering to their rehabilitation program can be challenging for healthcare providers  [kitzman2021physical]. Recent advances in technology have opened up the possibility of using machine learning algorithms to automatically assess patients’ performance of physical rehabilitation exercises. However, there are several challenges to overcome in order to achieve reliable and accurate assessment, including sensor placement, data variability, and patient heterogeneity.
Various technologies, including inertial measurement units (IMUs), computer vision, and machine learning algorithms, have been explored for automatically assessing physical rehabilitation exercises  [du2021assessing]. IMUs are small sensors attached to a patient’s body that measure movement and orientation, while computer vision techniques track joint positions and movements from video footage. Machine learning algorithms  [burns2018shoulder], like artificial neural networks, recognize exercise patterns and provide personalized feedback on patient performance using sensor data or video footage. While these approaches show promise in accurately identifying exercises and detecting errors, challenges remain in developing reliable and robust algorithms that can handle data variability and patient heterogeneity. More extensive validation studies in clinical settings are necessary to ensure the safety and effectiveness of these automated systems. Moreover, these approaches consider the exercise evaluation as a binary classification task, distinguishing exercises as either correct or incorrect. Nevertheless, these investigations were constrained in their capacity to furnish comprehensive feedback for individual exercise performances, as depicted in Figure 1.
Figure 1: Physical rehabilitation exercises process overview.
Deep learning approaches  [panwar2019rehab, zhang2020rehabilitation, liao2020deep] , including CNNs, RNNs, and GCNs, have received attention in the physical rehabilitation field for their ability to automatically assess patients’ exercise performance by analyzing sensor data or video footage. These models can recognize exercise patterns, detect errors, and provide feedback to patients, while also being able to handle patient heterogeneity and data variability. The use of Graph convolutional networks (GCNs), which operate on data represented as graphs, offers a unique advantage in capturing spatiotemporal characteristics of exercises and modeling complex interactions between body parts while respecting the natural topological structure of the human body skeleton.
Inspired by Graph Convolutional Networks (GCNs), we propose a novel Multi-Residual Spatio Temporal Graph Network (MR-STGN) that employs both information from positional and angular 3D skeletons to capture the spatial and temporal dependencies of human body parts. GCNs have shown promising results in modeling graph-structured data, and we leverage their benefits in our proposed model. The use of 3D skeletons allows us to capture the positional and angular information of human body parts, which is crucial in assessing patient actions during rehabilitation exercises. We employ a multi-residual framework to learn deeper representations and mitigate the problem of vanishing gradients in deep neural networks. Additionally, our proposed model utilizes a graph attention mechanism that selectively attends to relevant nodes and edges in the graph, improving the model’s discriminative power. Furthermore, our model provides clear guidance on which body parts or movements to focus on, leading to improved assessment quality. The real-time performance of the MR-STGAN model makes it suitable for healthcare applications requiring timely feedback, such as rehabilitation.
The main contributions of this paper are summarized as follows: (1) A novel Multi-Residual Spatio Temporal Graph Network (MRSTGN) that incorporates both angular and positional 3D skeletons for patient action assessment; (2) A multi-residual architecture that captures the spatio-temporal dynamics of patient actions by integrating information from multiple residual layers, each of which extracts features at different levels of abstraction; (3) An attention fusion mechanism that selectively combines the information from the angular and positional skeletons, allowing the network to adaptively weight the importance of different features and providing a more comprehensive and accurate assessment of patient action; (4)Guidance: The system provides clear guidance on which body parts or movements to focus on and has shown high performance on physical rehabilitation benshmarks; (5) High performance: The efficiency of the proposed model is shown through extensive experimentation on physical rehabilitation datasets UI-PRMD.
II Related Works
Rehabilitation exercises play a crucial role in the recovery process for patients with various health conditions. Evaluation of the quality and effectiveness of rehabilitation exercises is an essential task for healthcare professionals to ensure that patients receive optimal care. 3D skeleton-based approaches have emerged as a promising method for evaluating patient rehabilitation exercises due to their ability to capture the patient’s posture and movement accurately. There are several approaches for evaluating patient rehabilitation exercises based on 3D skeletons.
Discrete Movement Score (DMS) approaches
These approaches involve quantifying specific movements performed by patients during rehabilitation exercises and assigning a score based on the quality of the movement. One example of a DMS approach is the Fugl-Meyer assessment [gladstone2002fugl], which is commonly used to evaluate the motor function of stroke patients. The Fugl-Meyer assessment consists of a series of tasks, each of which is scored based on the quality of the patient’s movement. Other examples of DMS approaches include the Motor Assessment Scale, the Action Research Arm Test, and the Modified Ashworth Scale. DMS approaches have been shown to be reliable and valid for evaluating patient rehabilitation exercises. A study by Abreu et al.  [abreu2017assessment] found that the Fugl-Meyer assessment had high inter-rater reliability and good validity when used to evaluate stroke patients. Another work by Van der et al.  [van2001responsiveness] suggests that the Motor Assessment Scale was a reliable and valid measure of motor function in stroke patients. However, in general DMS approaches have numerous limitations including: subjectivity in scoring, leading to inter-rater variability; limited assessment scope, focusing on specific movements rather than overall functional abilities; time-consuming and resource-intensive process, requiring trained professionals; limited generalizability to populations beyond the specific target group.
Template-based approaches
Template-based approaches are used to evaluate the performance of rehabilitation exercises performed by patients by comparing them to pre-defined templates or models of correct movements. These approaches involve creating a template of a correct movement, which is then used as a reference point for evaluating the patient’s performance during rehabilitation exercises. One example of a template-based approach is the GAITRite system [bilney2003concurrent], which is used to evaluate gait performance in patients. Other examples of template-based approaches [henderson2007virtual] include the Virtual Peg Insertion Test and the Virtual Reality Rehabilitation System. A study by Menz et al. [menz2004reliability] propose the GAITRite system to measure of gait performance in patients with Parkinson’s disease. Another study by Subramanian et al.. [subramanian2013arm] found that the Virtual Reality Rehabilitation System was a valid and effective tool for upper limb rehabilitation in stroke patients. These approaches are limited by the lack of individualization, limited variability, and inflexibility to adaptations. They may struggle to capture complex movements accurately and have a narrow scope of assessment. These limitations highlight the need for considering alternative assessment methods to provide a comprehensive evaluation of patient progress and functional abilities during rehabilitation exercise
Deep Learning based approaches
Deep learning approaches have been increasingly used in recent years to evaluate patient rehabilitation exercises. These methods provide a powerful framework for analyzing complex, multi-modal data from different sources, such as motion sensors, video, and physiological signals, to provide accurate and reliable assessments of patient performance. One of the most popular deep learning models used for this task is the convolutional neural network (CNN). CNNs can learn discriminative features from raw sensor data and classify different types of exercises based on their patterns. For instance, researchers have used CNNs to classify different yoga poses using data from wearable motion sensors [anand2022yoga, wu2019yoga]. Another popular approach is to use graph convolutional networks (GCNs) to capture the spatial relationships between different parts of the body during exercise movements. GCNs can model the human body as a graph, where the nodes represent the body parts and the edges represent the connections between them. Researchers have used GCNs to predict the muscle activity of different body parts during squats [song2021secure] and to detect the correct execution of sit-to-stand exercises in patients with Parkinson’s disease [guo2020sparse]. The investigation of graph convolutional network (GCN) based approaches for rehabilitation holds great promise in advancing the field providing the ability to capture spatial relationships between body parts during exercise movements. This article proposes a novel GCN-based approach for accurately and in real time evaluation of rehabilitation exercises.
III Proposed approach
In this section, we introduce MR-STGAN, our proposed approach that incorporates a multi-residual spatio-temporal graph with attention fusion mechanisms.
III-A Problem Formulation
In this study, each exercise of rehabilitation is represented by an RGBD video denoted by $V_{i}={X_{t=1...T}}$, where $X_{t}$ is the $t^{ith}$ frame, and a ground-truth performance score $y_{i}\in[min_{score},max_{score}]$. To ensure the robustness of the data to changes in body sizes, motion rates, camera perspectives, and interference backgrounds, skeleton-based data encoding is chosen, where each skeleton has $N$ joints with C-dimensional coordinates and angles. The goal is to predict the score $\hat{y}_{j}$ and provide feedback to the patient to improve their exercise fluency. The proposed MR-STGAN model predicts the ground-truth performance score and highlights articulations that require improvement to provide feedback to the patient. The flowchart of our proposed approach is depicted in Figure 2. To ensure robustness to changes in body sizes, motion rates, camera perspectives, and interference backgrounds, we opted for skeleton-based data encoding over RGB image-based modality. Our encoding represents a sequence of skeletons, with each skeleton having $N$ joints, where each joint has C-dimensional coordinates. We encode these coordinates using a pose estimation approach or the sensor that captured the data. This encoding results in $V_{i}\in\mathbb{V}^{T\times N\times C}$ for each video and $X_{t}\in\mathbb{R}^{N\times C}$ for each frame. Our objective is to predict the score $\hat{y}_{j}$ for a given exercise, where each skeleton’s motion plays a crucial role. We also aim to provide feedback to the patient to improve the fluency of their exercise. To achieve this, we use a self-attention map $M_{j}\in\mathbb{R}^{T\times N\times N}$, which identifies the joints that need improvement and highlights them for the patient.
Figure 2: Flowchart of the proposed approach.
III-B Multi-Residual Spatio Temporal Graph Network (MR-STG)
[…]
$G(Z)=(Z.A).W$ (1)
where $W$ is a trainable weight matrix that is shared across all graphs to capture common features. For each kernel, $Z.A$ computes the weighted average of a node’s features with those of its neighboring nodes, which is multiplied by $(Z.A).W$, a weight matrix. This operation helps to extract spatial features from the non-linear structure of the skeletal sequence. Then, we use a Convolutional Gate Recurrent Unit (GRU) layer to compute a self-attention map which makes the adjacency matrix dynamic and is recomputed each time through the added layer. Compared to RNN and LSTM, ConvGRU is low time-consuming due to its simpler structure with gates in one unit. The benefits of using ConvGRU instead of other layers in terms of computation time and accuracy on metrics were evaluated in the experimental results. ConvGRU combines CNN and GRU, maintaining the spatial structure of the input sequence and facilitating the extraction of spatial-temporal features. For an action sequence ${x_{t=1...T}}$ with $T$ frames, forward propagation is performed as follows:
$z_{t}=\sigma(w_{zx}\otimes x_{t}+w_{zh}\otimes h_{t-1}+b_{z})$
$r_{t}=\sigma(w_{rx}\otimes x_{t}+w_{rh}\otimes h_{t-1}+b_{r})$
$o_{t}=tanh(w_{ox}\otimes x_{t}+w_{oh}\otimes(r_{t}\times f_{t-1})+b_{o})$
$h_{t}=z_{t}\times x_{t}+(1-z_{t})\times o_{t}$
Here, $\otimes$ and $\times$ denote the convolution operation and Hadamard product, respectively. $tanh$ and $\sigma$ represent tangent and sigmoid functions. $w_{x}$, $w_{h}$, and $b$ are corresponding weights and biases. $h_{t}$ denotes the hidden state for each time index $t=1..T$ ($h_{0}$ is set to 00) and is considered both the output and background information flowing into the network. The different gates in GRU are represented by $z_{t}$, $r_{t}$, and $o_{t}$. Finally, we inject an adjacency matrix $A$ by element-wise multiplication with the output of ConvGRU, and apply a normalization factor. The resulting output is then processed through a set of residual blocks, which consist of multiple residual layers. Finally, we average the output of the latter.
III-C Attention Fusion
The use of Attention fusion between spatio-temporal angular and positional features ($F_{positional}$ and $F_{angular}$) is a significant contribution to the field of patient action assessment. By combining these two types of features, the proposed method is able to provide a more comprehensive understanding of the patient’s movements during physical therapy exercises. To achieve this, the input tensor is split into two branches: one for the positional data and one for the angular data.
$C=[F_{positional},F_{angular}]$
The concatenated tensor $C$ is then passed through a dense layer with units=256 and activation=’relu’, followed by another dense layer with units=1 and activation=’softmax’. This results in a tensor of shape (batch, timesteps, 1) representing the attention weights for each frame.
$Att=softmax(relu(C*W1+b1)*W2+b2)$
where $W1$ and $W2$ are weight matrices, $b1$ and $b2$ are bias vectors, relu is the rectified linear unit activation function, * denotes element-wise multiplication and softmax function is used to ensure that the attention weights sum to one. The attention weights are then flattened using the Flatten() layer $F$ and repeated $N_{features}$ times along the second axis using the RepeatVector() layer.
$F=repeat(Att,N_{features})$
The tensor is then transposed using the Permute() layer to have a shape of (batch, $N_{features}$, timesteps) to allow element-wise multiplication with the concatenated tensor.
$D=C*F$
Finally, the attended features $D$ are pooled using global average pooling with the GlobalAveragePooling1D() layer to obtain a tensor of shape (batch, $N_{features}$), which represents the fused features for each example in the batch. This method combines the advantages of both angular and positional features to better capture the nuances of human motion. The angular features provide information on the orientation of the body segments, while the positional features provide information on the spatial location of the body segments. By combining these two types of features, the model can better understand the relationship between the different body segments and their movements. The attention mechanism further enhances the model’s ability to focus on the most informative parts of the input. The attention weights are learned during the training process and allow the model to selectively attend to certain parts of the input at different times. This means that the model can dynamically adjust its focus as the action progresses, which is particularly important for assessing complex actions. In addition, the use of attention fusion between the concatenated spatio-temporal angular and positional 3D skeletons features helps to address some of the limitations of previous methods. For example, previous methods [deb2022graph] have relied solely on either angular or positional features, which may have limited their ability to capture all aspects of human motion.
IV Experiments
This section details the evaluation of the proposed model MR-STGN. Comparative experiments were conducted to assess its performance, including the use of rehabilitation exercise datasets and evaluation metrics. Implementation details were discussed, followed by ablation studies that examined the impact of individual components. Lastly, the model was quantitatively compared to other state-of-the-art approaches.
IV-A Datasets and Metrics
Rehabilitation exercise dataset UI-PRMD was used in extensive experiments. The UI-PRMD dataset is publicly available and contains movements of common exercises performed by patients in physical rehab programs. Ten healthy individuals performed 10 repetitions of various physical therapy movements, captured using a Vicon optical tracker and a Microsoft Kinect sensor. The dataset includes full-body joint positions and angles, and its purpose is to serve as a foundation for mathematical modeling of therapy movements and establishing performance metrics to evaluate patients’ consistency in executing rehabilitation exercises. To evaluate and compare our approach with state-of-the-art methods, we present the evaluation of the mean absolute error (MAE) by comparing the predicted values with the actual performance scores on the validation set for a set of 10 exercises. The MAE was calculated for each exercise, which comprised a series of N sequences, using the following methodology:
$MAE=\frac{1}{N}\sum_{i=1}^{N}\|\hat{y_{i}}-y_{i}\|$
where $\hat{y_{i}}$ and $y_{i}$ denote the target variable and the prediction respectively.
Exercise Ours with both Only positional data (ours) Only angular data (ours) Deb et al. [deb2022graph] Du et al. [du2021assessing] Deep CNN [liao2020deep] Liao et al. [liao2020deep] Deep LSTM [liao2020deep] Co-occurence [li2018co] PA-LSTM [shahroudy2016ntu] Hierarchical LSTM [vakanski2018data] Two-stream CNN [simonyan2014two] Ex1 0.008 0.012 0.011 0.012 0.009 0.013 0.010 0.016 0.010 0.018 0.030 0.288 Ex2 0.010 0.014 0.011 0.011 0.020 0.029 0.028 0.049 0.029 0.044 0.077 0.223 Ex3 0.010 0.016 0.013 0.015 0.036 0.041 0.040 0.093 0.055 0.080 0.137 0.204 Ex4 0.008 0.011 0.010 0.010 0.014 0.016 0.011 0.016 0.013 0.023 0.035 0.360 Ex5 0.007 0.013 0.014 0.010 0.014 0.013 0.018 0.025 0.016 0.031 0.063 0.123 Ex6 0.010 0.015 0.018 0.017 0.020 0.023 0.017 0.021 0.018 0.034 0.046 0.211 Ex7 0.020 0.021 0.025 0.023 0.021 0.033 0.038 0.041 0.027 0.049 0.192 0.050 Ex8 0.020 0.026 0.021 0.024 0.022 0.029 0.023 0.046 0.024 0.050 0.072 0.043 Ex9 0.014 0.017 0.019 0.017 0.025 0.024 0.022 0.044 0.027 0.043 0.065 0.144 Ex10 0.015 0.026 0.028 0.025 0.026 0.036 0.041 0.052 0.046 0.077 0.160 0.110 Average 0,012 0,017 0,017 0,016 0,021 0,026 0,025 0,040 0,027 0,045 0,088 0.175
TABLE I: Results of our method in comparison with state-of-the-art approaches on the UI-PRMD dataset.
Phase Number of Videos Execution Time Deb et al. [deb2022graph] Train 373 58 hours Our model Train 373 29 min Deb et al. [deb2022graph] Test 100 13.87 seconds Our model Test 100 3.68 seconds
TABLE II: Computational time for Ex5 of UI-PRMD dataset for 1500 epochs.
IV-B Experimental Results
The MR-STGN model was implemented in Python 3.6 using the TensorFlow 2.x framework on a PC with an Intel Xeon Silver 4215R CPU, 32GB of RAM, and a GeForce GTX 3080 Ti 16GB RAM graphics card. The model was trained for 1500 epochs using the Adam optimizer and batch sizes of 4. The learning rate was set to $41e-$ 4. The choice of the regression loss function is critical to the performance of a model. Three different regression losses, namely MSE, Log-Cosh, and Huber loss, were used to evaluate the proposed model’s performance. The results show that Huber loss performs better for predicting assessment scores. Huber loss is advantageous as it improves the robustness of the model and provides a balance between MSE and MAE. It is also less sensitive to outliers, making the optimization process more stable and less prone to extreme values in the data. The best model was selected based on its performance on the validation set and was evaluated on the test set. To compare the model’s performance with recent works, 10 runs of the model were performed on the training and testing datasets. The performance measures MAE were saved from each run to ensure accuracy before averaging the results. we present the results obtained from various state-of-the-art approaches, such as the spatio-temporal neural network model, deep CNN model, and deep LSTM model as proposed in [liao2020deep]. Additionally, we include the co-occurrence model [li2018co], PA-LSTM [shahroudy2016ntu], two-stream CNN [simonyan2014two], and hierarchical LSTM [vakanski2018data]. Within our evaluation, we offer the findings derived from performing an ablation study on our model. This study involved using the same parameters with variations in data input, specifically focusing on positional data alone, angular data alone, and a combination of both.
Figure 3: Feedback visualization for different user profiles on UI-PRMD dataset: expert, not an expert with a good score, and not an expert with a low score. The attention Vector denotes the joint role vector (hot colors represent high values). Colored circles on the skeleton bodies allow the visualization of the attention vector and the role of body joints for different exercises.
Figure 4: Joints in the skeletal model of Kinect-recorded data from UI-PRMD dataset.
V Conclusion
In conclusion, this paper proposes a novel Multi-Residual Spatio Temporal Graph Network (MR-STGAN) that utilizes 3D skeletons and graph convolutional networks to assess patient actions during physical rehabilitation exercises. The proposed model employs a multi-residual architecture that captures the spatio-temporal dynamics of patient actions by integrating information from multiple residual layers. An attention fusion mechanism is utilized to selectively combine the information from the angular and positional skeletons, allowing the network to adaptively weight the importance of different features and provide more accurate and personalized feedback to patients. The performance of the proposed model was evaluated on physical rehabilitation dataset UI-PRMD, and it was found to outperform existing state-of-the-art models in terms of quantitative measures. Additionally, the proposed model provides clear guidance on which body parts or movements to focus on, and it has shown high performance on physical rehabilitation dataset. The real-time performance of the model makes it suitable for healthcare applications requiring timely feedback. In future work, the authors intend to explore the possibility of continuous assessment and develop a graphical interface to provide visual feedback to patients.